OpenStudy (anonymous):

2) When is it best to solve a system using elimination

OpenStudy (psymon):

When it's easy to make a variables or when its very hard to solve by substitution. For example: 2x - y = 8 -x + 3y = 4 This is a good system to use elimination on because its very easy to multiply the bottom equation by 2 and then add to eliminate the x's. 4x^2 + 6y^2 = 10 3x^2 - 3y^2 = 5 This is also a good candidiate for the elimnation method because it would be very hard to do by substitution.

OpenStudy (luigi0210):

Hello @Ammart98 Welcome Openstudy :)

OpenStudy (luigi0210):

*to

OpenStudy (anonymous):

Thank you

OpenStudy (psymon):

Mhm ^_^

OpenStudy (luigi0210):

._.

OpenStudy (luigi0210):

btw if you have any questions or concerns on the site feel free to PM me or anyone else with an A next to their name, or any mods that are enclosed in purple :) Hope you enjoy the site.

OpenStudy (luigi0210):

Thanks psymon ;D

OpenStudy (psymon):

haha, of course xD